### v4

@@ 3.4 Dataset Analysis
[…]
@@
To streamline the presentation, the JSON schemas for the dataset files are consolidated and displayed below in a structured format. These schemas define the structure and content of various dataset components.
text classification: {
biased_text: string,
dimension: string,
label: string
Figure 5: Schema for Text Classification.
Token classification: {
text: string,
Biased words: string
Figure 6: Schema for Token Classification.
Aspects: {
text: string,
Aspect: string
Figure 7: Schema for Aspects.
Stereotypes: {
stereotype_topic: string,
target_group: string,
template: string
Figure 8: Schema for Stereotype Prompts.
Demographics and sentiments: {
text: string,
label: string,
sentiment: string
Figure 9: Schema for Demographic Template.
Language Generation: {
biased_text: string,
debiased_text: string
Figure 10: Schema for Debiased Text.
3.5 Benchmarking
We evaluate the dataset by benchmarking state-of-the-art open-source models, including small language models (BERT-like) and LLMs (e.g., Llama, Mistral), for tasks such as text classification, token classification, and benign language generation. Additionally, we assess biases in these datasets across different demographics. Metrics such as accuracy, F1-score, fairness metrics (e.g., bias rate), and knowledge retention in language generation are used to evaluate model performance and robustness in mitigating bias.
4 Experimental Setup
@@ 4.2 Evaluation Schema
Models to evaluate
The baselines for classification tasks include smaller, encoder-only models from the BERT family (DistilBERT, BERT, RoBERTa), along with LLMs such as Mistral-7B-Instruct-v0.2 and Llama2-7B-chat that are tested in few-shot and fine-tuning setups. Few-shot approaches for LLMs are reported because they have demonstrated superior performance compared to zero-shot settings in our preliminary experiments. We utilize 80% of the 50,000 data samples for training (40,000 samples) and 20% for testing (10,000 samples) for the text classification task, which includes bias labels, toxicity, and sentiment. For token classification, we use the same train-test split on a dataset of 27,800 samples.
For the language generation task, only LLMs, such as Llama2-7B-chat and Mistral-7B-0.2-instruct, are used because it is a language generation task for which these decoder-only baselines are well-suited. These LLMs are utilized via prompts in two demonstrations. They were also fine-tuned on our 8.3k benign language generated dataset.
For the bias quantification tasks, both BERT-like models and LLMs such as Llama2-7B and Mistral-7B are utilized during inference, with two examples provided for demonstration. We used 29.85k data points: 352 for stereotypes and 520 samples for demographic alignment in the token prediction task.
For fine-tuning LLMs, we implemented Quantized Language Model Optimization with Low Rank Adapters (QLoRA) [73], a Parameter-Efficient Finetuning (PEFT) [74]. QLoRA is an optimization approach to minimize memory requirements during fine-tuning. Key QLoRA settings that we applied are the use of 4-bit quantization (Rank=64, alpha=16, dropout=0.2) to manage GPU memory limits. We used a batch size of 16 for training and 8 for evaluation for LLM fine-tuning. For smaller language models, our hyperparamters are different and batch sizes are larger, since they easily fit in given memory and compute usage. Detailed hyperparameters for each tasks are detailed in C.
@@
For text classification, we also use out-of-distribution datasets, including MBIB [64, 68] with 433,677 records for linguistic bias classification, Toxigen [25] with approximately 9,000 records for toxicity classification processed into binary labels using the official code [^1] https://github.com/microsoft/TOXIGEN, and the Financial News Polar Sentiment dataset with 4,840 sentences categorized by sentiment.
Evaluation Metrics
Metrics for Classification: Classification tasks, including sequence and token classification, are evaluated using the standard metrics of precision, recall, and accuracy. For BERT-like encoder-only (small language) models, which are probability-based, predicted probabilities determine these metrics by applying a 0.5 threshold to classify instances as positive or negative. LLMs like Llama2-7B-chat and Mistral-7B-instruct models generate text and provide confidence scores for their predictions, so we adjust the classification threshold for calculating these metrics (following practise in [75] for LLMs evaluation for classification).
Metrics for Bias Quantification Bias Rate: The Bias Rate is defined as the proportion of biased sentences generated by a model when completing sentences that include demographic placeholders. For instance, if 80 out of 100 completions are identified as biased based on references to attributes such as gender or race, the Bias Rate would be 80
Consider the template: “The [identity_mention] engineer is always late to meetings because she is juggling multiple family responsibilities and caregiving duties”, which is labeled as biased. If an LLM consistently completes the sentence as “The female engineer is always late to meetings because she is juggling multiple family responsibilities and caregiving duties”, this demonstrates a bias tendency in the LLM, specifically against the female gender.
Metrics for Language Generation We evaluate bias, toxicity, knowledge retention, faithfulness, and answer relevancy in LLM outputs both before and after applying prompting or fine-tuning techniques using our dataset. To conduct this evaluation, we use LLM-based scoring metrics and rely on GPT-4-Turbo as the evaluator. These metrics are taken from DeepEval [76].
The metrics are defined as follows: Bias and toxicity are measured as the percentage of outputs identified as biased or toxic, respectively, with lower percentages indicating better performance. Knowledge retention assesses how well the LLM preserves factual information from the original data. Faithfulness evaluates the accuracy and truthfulness of the generated content by comparing it against the original source. Answer relevancy determines whether the generated responses are relevant and appropriately aligned with the given input. For bias and toxicity, a lower score reflects better performance, while for knowledge retention, faithfulness, and answer relevancy, higher scores indicate improved results. Further details about the metrics can be found in C.4.
@@ 5 Results
In this section, we present results from evaluating the performance of different language models on the gold dataset [^2] Disclaimer: Some examples in this section are for demonstration and may not be found exactly in the data..
@@ 5.1 Text Classification for Bias, Toxicity, and Sentiment
Table 6: Classification Evaluation: Accuracy, precision, and recall metrics are reported for each classification category: Bias (‘0-Non-Bias’, ‘1-Bias’), Toxic (‘0-Non-Toxic’, ‘1-Toxic’), and Sentiment (‘0-Neutral’, ‘1-Positive’, ‘2-Negative’). Higher values are better and highlighted in bold.
@@
DistilBert-base (Fine-tuned) 0.85 0.81 / 0.90 0.91 / 0.79
Bert-large (Fine-tuned) 0.85 0.83 / 0.87 0.88 / 0.81
RoBERTa-large (Fine-tuned) 0.84 0.83 / 0.84 0.84 / 0.83
Llama2-7B (Few-shot) 0.61 0.59 / 0.65 0.76 / 0.45
Llama2-7B (Fine-tuned) 0.77 0.75 / 0.80 0.82 / 0.72
Mistral-7B-v0.2 (Few-shot) 0.53 0.56 / 0.48 0.66 / 0.38
Mistral-7B-v0.2 (Fine-tuned) 0.79 0.79 / 0.78 0.80 / 0.77
@@
DistilBert-base (Fine-tuned) 0.84 0.84 / 0.84 0.84 / 0.84
Bert-large (Fine-tuned) 0.85 0.85 / 0.84 0.84 / 0.85
RoBERTa-large (Fine-tuned) 0.85 0.87 / 0.82 0.81 / 0.88
Llama2-7B (Few-shot) 0.59 0.55 / 0.60 0.70 / 0.40
Llama2-7B (Fine-tuned) 0.75 0.73 / 0.78 0.80 / 0.70
Mistral-7B-v0.2 (Few-shot) 0.52 0.50 / 0.45 0.60 / 0.35
Mistral-7B-v0.2 (Fine-tuned) 0.77 0.76 / 0.77 0.78 / 0.75
@@
DistilBert-base (Fine-tuned) 0.87 0.96 / 0.79 / 0.88 0.84 / 0.94 / 0.83
Bert-large (Fine-tuned) 0.87 0.85 / 0.90 / 0.87 0.89 / 0.86 / 0.86
RoBERTa-large (Fine-tuned) 0.81 0.92 / 0.72 / 0.84 0.83 / 0.89 / 0.72
Llama2-7B (Few-shot) 0.58 0.50 / 0.55 / 0.55 0.65 / 0.40 / 0.45
Llama2-7B (Fine-tuned) 0.74 0.70 / 0.75 / 0.75 0.80 / 0.65 / 0.70
Mistral-7B-v0.2 (Few-shot) 0.50 0.45 / 0.40 / 0.45 0.55 / 0.30 / 0.35
Mistral-7B-v0.2 (Fine-tuned) 0.75 0.73 / 0.70 / 0.73 0.77 / 0.70 / 0.75
@@
Format:Each dataset entry follows the format text, label, with examples such as: “The movie was surprisingly good” (Positive), “Women are less capable in STEM fields compared to men” (Biased), and “You are worthless and nobody cares about you” (Toxic). The datasets for training and testing were sourced from BEAD Text Classification. For this study, we fine-tuned smaller BERT-like models and autoregressive LLMs, including Llama-2-7B-chat [77] and Mistral-7B-Instruct-v0.2 [3], using QLoRA [73]. These models were evaluated under both fine-tuned and few-shot settings. Unlike BERT-based models, which output probability scores, autoregressive LLMs yield categorical outputs. To ensure consistency, we used standard metrics like precision, recall, and accuracy, as described in [75]. Results: Table 6 illustrates that fine-tuned BERT-like models consistently outperformed autoregressive LLMs for bias classification, achieving approximately 85% accuracy. This finding underscores their efficiency and suitability for real-world applications, where computational resources might be limited.
Performance discrepancies between smaller models and LLMs (which have billions of parameters) likely stem from differences in architectural design, dataset sizes, and task-specific nuances. Further investigation into these factors, as well as an analysis of out-of-distribution datasets such as MBIB [64], Toxigen [25], and financial news [78], reinforces these observations. Results for these datasets, detailed in Figure11, align with the conclusions presented here.
We also evaluate the performance of these models on out-of-distribution datasets, including MBIB [64, 68] for bias classification, Toxigen [25] for toxicity classification, and the Financial News Polar Sentiment dataset for sentiment analysis, as shown in Table 11. Same train-val-test spltis are used for all datasets. The results indicate that these models perform better with the BEADs dataset, highlighting its suitability for fine-tuning and evaluation tasks.
@@ 5.2 Token Classification Task
Format: Dataset entries are structured as: sentence, array of biased labeled tokens (e.g., “Women are overreacting bossy managers.”, [‘overreacting’, ‘bossy managers’]). This dataset is available as the Token Classification Dataset.
For this task, we fine-tuned smaller models such as BiLSTM-CRF, BERT-like models, and LLMs, including Llama2-7B-chat and Mistral-7B-instruct. The models were evaluated on their ability to classify tokens as ‘Bias’ or ‘O’ (outside bias). We formatted the data following CoNLL-2003 standards [79] for NER, presenting it as the CoNLL Bias Dataset. The standard ‘B-BIAS’ and ‘I-BIAS’ tags were merged into ‘Bias’, while the ‘O’ tag was retained as ‘Non-Bias’ for evaluation purposes.
Table 7: Token Classification Evaluation: Metrics include accuracy, precision, and recall for ’Non-Bias’ (0) and ’Bias’ (1) entities. Higher ($\uparrow$) values indicate better performance and are highlighted in bold.
@@
DistilBert-base (Fine-tuned) 0.80 0.67 / 0.84 0.62 / 0.88
Bert-large (Fine-tuned) 0.81 0.61 / 0.87 0.65 / 0.85
RoBERTa-large (Fine-tuned) 0.82 0.72 / 0.89 0.70 / 0.89
BiLSTM-CRF (Fine-tuned) 0.60 0.42 / 0.45 0.42 / 0.55
Llama2-7B (Few-shot) 0.74 0.61 / 0.73 0.57 / 0.76
Llama2-7B (Fine-tuned) 0.76 0.64 / 0.76 0.60 / 0.79
Mistral-7B-v0.2 (Few-shot) 0.78 0.62 / 0.78 0.63 / 0.81
Mistral-7B-v0.2 (Fine-tuned) 0.80 0.65 / 0.80 0.67 / 0.83
Results: Table 7 highlights that smaller fine-tuned BERT-like models achieved superior performance across all metrics, effectively classifying both ‘Non-Bias’ and ‘Bias’ tokens. RoBERTa-large, in particular, emerged as the best-performing model, with the highest accuracy, precision, and recall.
Few-shot models, such as Llama2-7B and Mistral-7B, demonstrated moderate performance, with fine-tuning significantly improving their token classification capabilities. Notably, fine-tuned LLMs consistently outperformed their few-shot counterparts.
These findings align with trends observed in other tasks: smaller models often outperform larger LLMs in specific, well-defined tasks like token classification. This disparity may be due to smaller models’ efficiency in processing and adapting to limited data, faster convergence during fine-tuning, or reduced risk of overfitting compared to LLMs with significantly larger parameter spaces.

### v5

@@ 3.4 Dataset Analysis
The hiring committee is right to decide that women tend to be too emotional for leadership roles and that men are naturally more decisive.
@@
To facilitate clearer interpretation of our 1,876 fine-grained aspect labels, we introduce the following taxonomy that groups them into ten high-level categories. Table 7 summarizes each category, its definition, and representative examples.
Table 7: Aspect taxonomy: grouping fine-grained labels into ten high-level categories.
Category Description Examples
Personal Traits and demographics ability, identity, age, gender, accent
Emotional Affective states acceptance, anxiety, happiness, anger
Social Interpersonal relations abuse, support, discrimination, friendship
Activities Actions and events accident, achievement, acquisition, action
Resources Physical/digital access access, funding, equipment, opportunity
Abstract Intangible concepts accountability, fairness, accuracy, justice
Temporal References to time or location time, date, location, absence
Entities Concrete items or organizations account, organization, product, event
Health Physical and mental well-being health, safety, risk, injury
Technical Systems, processes, or methods performance, functionality, method, metric
To enable analyses of model behavior across different population groups, each record in our dataset is annotated with six demographic dimensions. Table 8 summarizes these dimensions, their definitions, and the discrete labels we employ.
Table 8: Demographic dimensions and their corresponding labels.
Dimension Description Labels
Age Group Broad age brackets Under 18; 18–34; 35–54; 55+
Gender Self-reported gender identity Female; Male; Non-binary; Prefer not to say
Ethnicity Self-reported racial/ethnic background White; Black; Asian; Hispanic; Other
Education Level Highest formal educational attainment High school or less; Some college; Bachelor’s; Graduate
Employment Status Current work status Employed; Unemployed; Student; Retired
Geographic Region World region of residence North America; Europe; Asia-Pacific; Latin America; Africa; Other
1{
2 "text_classification": {
3 "biased_text": "string",
4 "dimension": "string",
5 "label": "string"
6 }
7}
(a) Text Classification
1{
2 "token_classification": {
3 "text": "string",
4 "biased_words": "string[]"
5 }
6}
(b) Token Classification
1{
2 "aspects": {
3 "text": "string",
4 "aspect": "string"
5 }
6}
(c) Aspects
1{
2 "stereotypes": {
3 "topic": "string",
4 "target_group": "string",
5 "template": "string"
6 }
7}
(d) Stereotype Prompts
1{
2 "demographics_sentiment": {
3 "text": "string",
4 "label": "string",
5 "sentiment": "string"
6 }
7}
(e) Demographic Template
1{
2 "language_generation": {
3 "biased_text": "string",
4 "debiased_text": "string"
5 }
6}
(f) Debiased Text
Figure 7: JSON schemas defining each dataset file’s structure.
@@
Figure 7 presents the six JSON schema definitions corresponding to each component of the BEADs dataset. These schemas illustrate the structure and annotation format used across the six tasks, ranging from classification to generation.
4 Benchmarking Setup
@@ 4.2 Evaluation Schema
4.2.1 Models Evaluated
Our evaluation covers both small encoder-only models, such as DistilBERT, BERT, and RoBERTa, and decoder-only large language models (LLMs), specifically Mistral-7B-Instruct-v0.2 and Llama2-7B-chat. For classification tasks, both categories are considered. For generation tasks, only decoder-style LLMs are used, as they are better suited to generative outputs. In addition to fine-tuning, we explore prompting-based evaluation for all LLMs, leveraging few-shot demonstrations to assess their performance in instruction-following and bias-sensitive contexts. Preliminary trials showed that few-shot prompting notably outperformed zero-shot setups, which motivated its use throughout our evaluation.
4.2.2 Training and Evaluation Settings
For the text classification task (bias, toxicity, and sentiment), we use 50,000 examples split into 40,000 training and 10,000 testing samples (80/20 split). For token classification, the same proportion is applied to a dataset of 27,800 annotated examples. For the language generation task, only LLMs are employed, prompted with two-shot examples and optionally fine-tuned on our curated dataset of 8,300 benign language rewrites.
For bias quantification tasks, both BERT-style models and LLMs are evaluated in inference-only settings using two illustrative prompts. This component includes 29,850 samples: 352 for stereotype probing and 520 for demographic alignment detection via token prediction.
4.2.3 Fine-Tuning Strategy
For LLM fine-tuning, we apply Quantized Language Model Optimization with Low-Rank Adapters (QLoRA) [72], which enables parameter-efficient finetuning (PEFT) [73] on limited hardware. We use 4-bit quantization with rank=64, alpha=16, and dropout=0.2 to manage memory constraints. LLMs are trained with a batch size of 16 and evaluated with a batch size of 8. Smaller encoder-only models are fine-tuned with larger batch sizes due to their lower memory footprint. Full details of model-specific hyperparameters are provided in  Appendix C.
@@ 4.2.3 Fine-Tuning Strategy
To assess text classification generalization, we test our models on three out-of-distribution datasets. The MBIB dataset [37, 68] contains 433,677 records for linguistic bias classification. Toxigen [25], with approximately 9,000 samples, is used for toxicity classification after binarizing labels using the official codebase [^1] https://github.com/microsoft/TOXIGEN. We also use the Financial News Polar Sentiment dataset, which comprises 4,840 labeled sentences categorized by sentiment polarity.
4.2.4 Evaluation Metrics
Classification tasks, including both sequence-level and token-level classification, are evaluated using standard metrics: precision, recall, and accuracy. For BERT-style models, a threshold of 0.5 is applied to model-predicted probabilities. LLMs, by contrast, generate textual responses and self-reported confidence scores, requiring adjusted classification thresholds. We follow best practices outlined in Inan et al. [74] for calibrating these evaluations.
Bias quantification uses the Bias Rate, defined as the proportion of model completions identified as biased when prompted with identity-sensitive templates. For instance, completions such as “The female engineer is always late…” from prompts containing placeholders like [identity_mention] are flagged for gender bias. A higher Bias Rate indicates a stronger tendency toward demographic stereotyping.
For the language generation task, we measure five attributes: bias, toxicity, knowledge retention, faithfulness, and answer relevancy. These are computed using GPT-4 as an evaluator via the DeepEval framework [75]. Bias and toxicity are reported as the proportion of biased or toxic outputs, where lower values reflect improved performance. Knowledge retention assesses how well factual content is preserved from the input. Faithfulness measures the consistency of generated content with reference information. Answer relevancy evaluates alignment between the response and the original question. For the last three metrics, higher scores indicate better performance. Further details are provided in  Appendix C.4.
@@ 5 Results
In this section, we present results from evaluating the performance of different language models on the dataset [^2] Disclaimer: Some examples in this section are for demonstration and may not be found exactly in the data..
@@ 5.1 Text Classification for Bias, Toxicity, and Sentiment
Table 9: Classification metrics (Accuracy, Precision, Recall) across three tasks: Bias, Toxicity, and Sentiment. Higher values are better and highlighted in bold. FS stands for few-shot (2) prompting and FT stands for fine-tuning.
@@
DistilBERT-base (FT) 0.85 0.81 / 0.90 0.91 / 0.79
BERT-large (FT) 0.85 0.83 / 0.87 0.88 / 0.81
RoBERTa-large (FT) 0.84 0.83 / 0.84 0.84 / 0.83
Llama2-7B (FS) 0.61 0.59 / 0.65 0.76 / 0.45
Llama2-7B (FT) 0.77 0.75 / 0.80 0.82 / 0.72
Mistral-7B-v0.2 (FS) 0.53 0.56 / 0.48 0.66 / 0.38
Mistral-7B-v0.2 (FT) 0.79 0.79 / 0.78 0.80 / 0.77
@@
DistilBERT-base (FT) 0.84 0.84 / 0.84 0.84 / 0.84
BERT-large (FT) 0.85 0.85 / 0.84 0.84 / 0.85
RoBERTa-large (FT) 0.85 0.87 / 0.82 0.81 / 0.88
Llama2-7B (FS) 0.59 0.55 / 0.60 0.70 / 0.40
Llama2-7B (FT) 0.75 0.73 / 0.78 0.80 / 0.70
Mistral-7B-v0.2 (FS) 0.52 0.50 / 0.45 0.60 / 0.35
Mistral-7B-v0.2 (FT) 0.77 0.76 / 0.77 0.78 / 0.75
@@
DistilBERT-base (FT) 0.87 0.96 / 0.79 / 0.88 0.84 / 0.94 / 0.83
BERT-large (FT) 0.87 0.85 / 0.90 / 0.87 0.89 / 0.86 / 0.86
RoBERTa-large (FT) 0.81 0.92 / 0.72 / 0.84 0.83 / 0.89 / 0.72
Llama2-7B (FS) 0.58 0.50 / 0.55 / 0.55 0.65 / 0.40 / 0.45
Llama2-7B (FT) 0.74 0.70 / 0.75 / 0.75 0.80 / 0.65 / 0.70
Mistral-7B-v0.2 (FS) 0.50 0.45 / 0.40 / 0.45 0.55 / 0.30 / 0.35
Mistral-7B-v0.2 (FT) 0.75 0.73 / 0.70 / 0.73 0.77 / 0.70 / 0.75
@@
Format: Each dataset entry follows the format text, label, with examples such as: “The movie was surprisingly good” (Positive), “Women are less capable in STEM fields compared to men” (Biased), and “You are worthless and nobody cares about you” (Toxic). The datasets for training and testing were sourced from BEAD Text Classification.
For this study, we fine-tuned smaller BERT-like models and autoregressive LLMs, including Llama-2-7B-chat [76] and Mistral-7B-Instruct-v0.2 [3], using QLoRA [72]. These models were evaluated under both fine-tuned and few-shot settings. Unlike BERT-based models, which output probability scores, autoregressive LLMs yield categorical outputs. To ensure consistency, we used standard metrics like precision, recall, and accuracy, as described in [74].
Results: Table 9 illustrates that fine-tuned BERT-like models consistently outperformed autoregressive LLMs for bias classification, achieving approximately 85% accuracy. This finding underscores their efficiency and suitability for real-world applications, where computational resources might be limited. Performance discrepancies between smaller models and LLMs (which have billions of parameters) likely stem from differences in architectural design, dataset sizes, and task-specific nuances.
We also evaluate the performance of these models on out-of-distribution datasets, including MBIB [37, 68] for bias classification, Toxigen [25] for toxicity classification, and the Financial News Polar Sentiment dataset for sentiment analysis, as shown in Figure 8. Same train-val-test spltis are used for all datasets. The results indicate that these models perform better with the BEADs dataset, highlighting its suitability for fine-tuning and evaluation tasks.
@@ 5.2 Token Classification Task
Format: Dataset entries are structured as: sentence, array of biased labeled tokens (e.g., “Women are overreacting bossy managers.", [‘overreacting’, ‘bossy managers’]). This dataset is available as the Token Classification Dataset.
For this task, we fine-tuned smaller models such as BiLSTM-CRF, BERT-like models, and LLMs, including Llama2-7B-chat and Mistral-7B-instruct-v0.2. The models were evaluated on their ability to classify tokens as ‘Bias’ or ‘O’ (outside bias). We formatted the data following CoNLL-2003 standards [77] for NER, presenting it as the CoNLL Bias Dataset. The standard ‘B-BIAS’ and ‘I-BIAS’ tags were merged into ‘Bias’, while the ‘O’ tag was retained as ‘Non-Bias’ for evaluation purposes.
Table 10: Token Classification Evaluation: Metrics include accuracy, recall, and precision for ’Non-Bias’ (0) and ’Bias’ (1) entities. Higher ($\uparrow$) values indicate better performance and are highlighted in bold.
@@
DistilBERT-base (FT) 0.80 0.67 / 0.84 0.62 / 0.88
BERT-large (FT) 0.81 0.61 / 0.87 0.65 / 0.85
RoBERTa-large (FT) 0.82 0.72 / 0.89 0.70 / 0.89
BiLSTM-CRF (FT) 0.60 0.42 / 0.45 0.42 / 0.55
Llama2-7B (FS) 0.74 0.61 / 0.73 0.57 / 0.76
Llama2-7B (FT) 0.76 0.64 / 0.76 0.60 / 0.79
Mistral-7B-v0.2 (FS) 0.78 0.62 / 0.78 0.63 / 0.81
Mistral-7B-v0.2 (FT) 0.80 0.65 / 0.80 0.67 / 0.83
Results: Table 10 highlights that smaller fine-tuned BERT-like models achieved superior performance across all metrics, effectively classifying both ‘Non-Bias’ and ‘Bias’ tokens. RoBERTa-large, in particular, emerged as the best-performing model, with the highest accuracy, precision, and recall. Few-shot models, such as Llama2-7B-chat and Mistral-7B-Instruct-v0.2, demonstrated moderate performance, with fine-tuning significantly improving their token classification capabilities. Notably, fine-tuned LLMs consistently outperformed their few-shot counterparts.
These findings align with trends observed in text classification task also where smaller models often outperform larger LLMs in specifi. This disparity may be due to smaller models’ efficiency in processing and adapting to limited data, faster convergence during fine-tuning, or reduced risk of overfitting compared to LLMs with significantly larger parameter spaces.